Consider a normal distribution curve where the middle 20 % of the area under the curve lies above the interval ( 7 , 20 ). Use this information to find the mean, ? , and the standard deviation, ?, of the distribution.
Answer #1

From standard normal table:

P(-0.253 < z < 0.253) = 0.20

Hence,

(7 - )/ = -0.253

And

(20 - )/ = 0.253

On solving above two equations, we will have:

7 + 0.253 = 20 - 0.253

= 25.69

= 7 + 0.253*25.69 = 13.5